OpenStudy (abmon98):

particles A and B are moving along horizontal straight lines which lie in the west east direction. At a certain time, A passes a point P moving to the east with speed 5ms^-1 and constant acceleration 1ms^-2. Simultaneously, B passes a point Q east of P with speed 4ms^-1 towards the west but with constant acceleration 2ms^-2 towards the east. Given that the distance PQ is 28m, Calculate a) The distance of A from P when B reverse direction b) The time taken after the start for A to overtake c) The further time taken for B to overtake A

OpenStudy (lastdaywork):

Do you know the concept of relative motion ?

OpenStudy (abmon98):

v=u+at v=0 as B has reversed so 0=4-2t so t=2 seconds s=5(2)+1/2(2)^2=12 –

OpenStudy (abmon98):

thats the distance A took while particle B reversed in 2 seconds

OpenStudy (lastdaywork):

It's correct.

OpenStudy (lastdaywork):

Now try part (b)